Circular No. 5687

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993a) Jean Mueller reports her discovery of a comet on a plate taken during

the course of the second Palomar Sky Survey with the 1.2-m Oschin Schmidt Telescope, and provides the following positions from trails. The object is condensed with a prominent coma and a faint tail to the south.

1993 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 2.40833  9^{h}46^{m}45.^{s}34  +47^{o}02′27′′.8  15.5
2.45347  9 46 42.62  +47 03 05.2
3.52708  9 45 28.25  +47 20 23.7
3.54792  9 45 27.11  +47 20 37.7

SUPERNOVA 1992bo IN ESO 352-G57 M. Hamuy, Cerro Tololo Interamerican Observatory, reports that a

spectrogram (range 320-750 nm) of SN 1992bo (cf. IAUC 5682), obtained on 1992 Dec. 31 UT with the CTIO 1.5-m telescope, shows the characteristic Si II 635.5-nm feature of type-Ia supernovae. The minimum of the Si II absorption yields an expansion velocity of ∼ 10 500 km/s with respect to the galaxy’s rest frame. This spectrum shows resemblance to that of the type-Ia SN 1992A taken 1 day before maximum light. The spectrum of the parent galaxy’s nucleus yields a redshift of z = 0.019. Spectrophotometry yields synthetic magnitudes B = 15.9, V = 15.8 (± 0.1).

M. Della Valle, European Southern Observatory; and K. Reinsch, Universität Sternwarte, Göttingen, report that inspection of CCD spectrograms (range 380-960 nm, resolution ∼ 2 nm), obtained on Dec. 31.1 UT with the ESO/MPI 2.2-m telescope (+ EFOSC II) at La Silla, show SN 1992bo to be a type-Ia supernova at maximum light. The spectrum is dominated by strong absorptions of Si II (635.5, 597.2, 564.0, 545.4, and 412.9 nm), Fe II (501.8 nm), and Mg II (448.1 nm).

SUPERNOVA 1992bp IN ANONYMOUS GALAXY Hamuy also reports that a spectrogram of SN 1992bp (cf. IAUC 5682)

obtained on Dec. 31 UT shows the characteristic Si II 635.5-nm feature of type-Ia supernovae. The minimum of the Si II absorption yields an expansion velocity of ∼ 10 800 km/s with respect to the galaxy’s rest frame. This spectrum shows resemblance to that of the type-Ia SN 1992A taken seven days after maximum light; the parent galaxy has z = 0.080.

1993 January 4 Daniel W. E. Green